OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). cos x = sin x

OpenStudy (abb0t):

Set it equal to \[\tan(x) = 0\] on the given interval...

OpenStudy (abb0t):

\[\frac{ \sin(x) }{ \cos(x) } = \tan(x)\]

OpenStudy (anonymous):

tan(x)=1

OpenStudy (anonymous):

x=pi/4

OpenStudy (anonymous):

pi/4+pi=5pi/4